Formulate the situation as a system of two linear equations in two varlahles. Re sure to state clearly the meaning of yaur x- and y-ariahles. Solve the system by the ellmination method. Re sure to state your final answer In terms of the original question. A jar contains 80 nickels and dines worth $6.80. How many of each kind of coin are in the jer? nickels dimes Show My Work (optional
Answer #1

Let x be the number of nickels and y be the number of dimes
x+y=80 .....1)
and x*0.05+y*0.10=6.80 .....2)

From 1),
y=80-x
Substituting the value of y in 2), we get
x*0.05+(80-x)*0.10=6.80
=>x=24
Hence, y=56

x=24 nickels
y=56 dimes